OpenStudy (anonymous):

if c(x) = 4x – 2 and f(x) = x2 + 5x what is (c*d)(x)

OpenStudy (danjs):

\[C(x) *f(x) = (C*f)(x)\]

OpenStudy (danjs):

(4x-2)* (x^2+5x) = ...

OpenStudy (anonymous):

4x3+18x2−10x

OpenStudy (anonymous):

tyty

OpenStudy (danjs):

welcome